The electrolytic manufacture of chlorine gas from a sodium chloride solution is carried out by the following reaction: 2 NaCl + 2 H2O → 2 NaOH + H2 + Cl2 How many kilograms of Cly can be produced from 10 m3 of brine solution containing 5% by weight of NaCl? The specific gravity of the solution relative to that of water at 4°C is 1.07. ** Assuming complete converson of NaCl Atomic mass: Na: 23 Cl: 35.5 0:16 H: 1


1
Expert's answer
2021-06-10T03:35:09-0400

Density = m / V

m = density x V

Density = 1.07 kg/m3

Mass of 10 m3 NaCl solution = 1.07 x 10 = 10.7 kg

w% = (100 x m(NaCl)) / msolution

m (NaCl) = (w% x msolution) / 100

m (NaCl) = (5 x 10.7) / 100 = 0.535 kg

According to the equation, n (Cl2) = 1/2 n (NaCl)

n = m / M

m = n x M

M (NaCl) = 58 g/mol

M (Cl2) = 71 g/mol

n (NaCl) = 0.535 / 58 = 0.01 mol

n (Cl2) = 1/2 n (NaCl) = 0.005 mol

m (Cl2) = 0.005 x 71 = 0.33 kg